A glass container ( k=1.05 W/m-K) with a surface area 0.3m2 is filled with hot water at 100oC and was let to cool in a room with a temperature of 27oC. If the walls of the container are 0.8cm thick, how much heat is transferred from inside the glass to the room per second?


1
Expert's answer
2022-01-18T09:36:46-0500

Given:

"k=1.05\:\rm W/m\cdot K"

"A=0.3\:\rm m^2"

"t_1=100\:^\circ\rm C"

"t_2=27\:^\circ\rm C"

"\Delta h=0.008\:\rm m"


The Fourier law of heat conductivity says

"\frac{dQ}{dt}=k\frac{\Delta t}{\Delta h}A"

"\frac{dQ}{dt}=1.05*\frac{100-27}{0.008}*0.3=2874\:\rm W"